OpenStudy (highschoolmom2010):

An escalator in the subway station has a vertical rise of 195 ft 9.5 in., and rises at an angle of 10.4 degrees. How long is the escalator? Round to the nearest foot.

OpenStudy (highschoolmom2010):



OpenStudy (highschoolmom2010):

we know hyp=x opposite =2349.5

ganeshie8 (ganeshie8):

keep going, you have it !

OpenStudy (highschoolmom2010):

SOH CAH COA so we use sin

ganeshie8 (ganeshie8):

yes !

OpenStudy (highschoolmom2010):

\[(x)*\sin10.4=\frac{ 2349.5 }{ x }*(x)\] \[\frac{ xsin10.4 }{\sin 10.4 }=\frac{ 2349.5}{\sin 10.4 }\] \[x=\frac{ 2349.5 }{ \sin 10.4}\] \[x \approx 1085'\]

OpenStudy (highschoolmom2010):

that doesnt look right

ganeshie8 (ganeshie8):

1085 foot is correct ! good job !!

ganeshie8 (ganeshie8):

*feet

OpenStudy (highschoolmom2010):

hooray t just didnt look right

ganeshie8 (ganeshie8):

you're becoming extremely effecient at trig rations ;) good :D

OpenStudy (jdoe0001):

ahemm, wasn't 2349.5 in inches?

OpenStudy (highschoolmom2010):

i converted it back to feet

OpenStudy (highschoolmom2010):

bc the question wanted it in feet

OpenStudy (jdoe0001):

I see, you have a good footing